a body a body of mass 2 kg is placed on a rough inclined plane whose angle of inclination of 30-degree. find the resultant force acting down the slope to the acceleration. If a frictional force of 5 Newton acts between the body and the inclined surface.
1
Expert's answer
2020-04-13T10:04:50-0400


The resultant force acting down the slope is projection of force of gravity minus frictional force: "F = m g \sin \theta - F_f".

Given that "\theta = 30^{\circ}", "m = 2 kg" and "F_f = 5 N", the resultant force is "F = 2 kg \cdot 9.81 \frac{m}{s^2} \cdot 0.5 - 5 N = 4.81 N".